# Boltzmann priors for Implicit Transfer Operators

### v1

@@ Abstract
Accurate prediction of thermodynamic properties is essential in drug discovery and materials science. Molecular dynamics (MD) simulations provide a principled approach to this task, yet they typically rely on prohibitively long sequential simulations. Implicit Transfer Operator (ITO) Learning offers a promising approach to address this limitation by enabling stable simulation with time steps orders of magnitude larger than MD. However, to train ITOs, we need extensive, unbiased MD data, limiting the scope of this framework. Here, we introduce Boltzmann Priors for ITO (BoPITO) to enhance ITO learning in two ways. First, BoPITO enables more efficient data generation, and second, it embeds inductive biases for long-term dynamical behavior, simultaneously improving sample efficiency by one order of magnitude and guaranteeing asymptotically unbiased equilibrium statistics. Further, we showcase the use of BoPITO in a new tunable sampling protocol interpolating ITO models trained on off-equilibrium simulation data and an unbiased equilibrium distribution to solve inverse problems in molecular science.
@@ 1 Introduction
Here, we introduce Boltzmann Priors for Implicit Transfer Operator (BoPITO) learning (Figure 1). BoPITO leverages pre-trained Boltzmann Generators (BG) (Noé et al., 2019; Viguera Diez et al., 2024; Klein & Noé, 2024; Köhler et al., 2023; 2020; Midgley et al., 2024; 2023) to enable data-efficient training of ITO models, leading to one order of magnitude reduction in the simulation data needed for training. As the BG encodes the invariant measure or Boltzmann distribution of the dynamics encoded by the transition density, BoPITO, by construction, guarantees asymptotically unbiased equilibrium statistics. In this way, BoPITO can combine off-equilibrium data and biased data encoded into a BG prior to train ITO models that predict MD across multiple time-scales. Using BoPITO we introduce a new sampling strategy to recover approximate dynamics from biased off-equilibrium data, a BG prior, and unbiased time-correlation data, providing a new method for inverse problems for molecular systems.
@@ 2.1 Molecular dynamics and observables
Unfortunately, stable MD simulations rely on using time discretizations, $\tau$, on the order of $10^{-15}\,\mathrm{s}$, whereas properties such as protein folding or ligand unbinding occur on much longer time scales, $\sim 10^{-3}-10^{-1}\,\mathrm{s}$. Due to the sequential nature of these simulations, an impractical number of simulation steps are necessary to predict these properties in an unbiased fashion. As a result, most MD data will be ‘off-equilibrium’, e.g. trajectories exploring one or a few of the modes of $\mu(\mathbf{x})$.
@@ 2.2 Boltzmann Generators
Boltzmann Generators (BG) (Noé et al., 2019) are a generative machine learning approach to draw samples i.i.d. from the Boltzmann distribution, $\mu(\mathbf{x})$, of physical many-body systems (eq. 1). BGs learn a diffeomorphic map, $\mathcal{F}_{\bm{\theta}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$, from a latent space, equipped with a simple base density $p(\mathbf{z})$, to configurational space of a physical system, such that the push-forward density $\hat{p}(\mathbf{x})=\mathcal{F}_{\bm{\theta}}\#p(\mathbf{z})$ closely approximates the Boltzmann distribution (eq. 1). In practice, $\mathcal{F}_{\bm{\theta}}$ is implemented using an invertible neural network architecture with tractable Jacobian determinants to enable efficient sampling and exact sample likelihood computation (Chen et al., 2018; Papamakarios et al., 2021). BGs are trained either with approximately equilibrated simulation data or with biased simulations, from i.e. enhanced sampling, by employing appropriate reweighting (Ferrenberg & Swendsen, 1989; Shirts & Chodera, 2008). Unbiased samples can then be generated by importance re-sampling. Similarly, unbiased expectations can be computed using the importance weights $w(\mathbf{x})=e^{-\beta U(\mathbf{x})}/\hat{p}(\mathbf{x})$. As such, BGs are surrogates of the target Boltzmann distribution, but do not model time-correlation statistics.
@@ 2.3 Transfer operators
where $\lambda_{i}(\tau)=\exp(-\tau\kappa_{i})$ are the eigenvalues, $\langle f|g\rangle=\int_{\Omega}f(\mathbf{x})g(\mathbf{x})\,\mathrm{d}\mathbf{x}$ and $\psi_{i}$ and $\phi_{i}$ are the corresponding right and left eigenfunctions, respectively. The eigenvalues $\lambda_{i}(\tau)$ depend on $\tau$, and $\kappa_{i}$ are characteristic relaxation rates. When the dynamics are reversible, all eigenvalues $\lambda_{i}$ are real-valued and lie in the interval $-1<\lambda_{i}\leq 1$. Notably, there is one eigenvalue $\lambda_{1}=1$ and its associated eigenfunction corresponds to the stationary distribution $\mu(\mathbf{x})$. Moreover, the left and right eigenfunction are related by a factor of the stationary density, $\phi_{i}(\mathbf{x})=\mu(\mathbf{x})\psi_{i}(\mathbf{x})$, and $\psi_{1}(\mathbf{x})=\mathbf{1}$.
@@ 2.4 Diffusion models
$\mathrm{d}\mathbf{x}^{t}=f(\mathbf{x}^{t},t)\,\mathrm{d}t+g(t)\,\mathrm{d}W,$ (6)
where ($0<t<T$) is the diffusion time, $f$ and $g$ are chosen functions, and $\mathrm{d}W$ is a Wiener process. To generate samples from the data distribution, $p(x^{0})$, we can sample from $p(\mathbf{x}^{T})$ and solve the backward diffusion process (‘denoising process’) (Anderson, 1982),
$\mathrm{d}\mathbf{x}^{t}=\left[f(\mathbf{x}^{t},t)-g^{2}(t)\nabla_{\mathbf{x}^{t}}\log p(\mathbf{x}^{t}|t)\right]\mathrm{d}t+g(t)\,\mathrm{d}W.$ (7)
This backward process is approximated using a deep neural network $\nabla_{\mathbf{x}^{t}}\log p_{\bm{\theta}}(\mathbf{x}^{t}|t)=s_{\bm{\theta}}(\mathbf{x}^{t},t)$. The learned backward SDE can be recast into a probability flow ODE (Maoutsa et al., 2020; Song et al., 2021), which in turn can be interpreted as a continuous normalizing flow (Chen et al., 2018) which facilitates efficient sampling and enables sample likelihood evaluation.
@@ 3 Boltzmann Priors for Implicit Transfer Operator Learning
BoPITO uses pre-trained models of the equilibrium distribution to improve ITO learning in three ways. First, it helps ensure broad sampling of $\Omega$ proportional to the equilibrium distribution for subsequent MD simulations yielding information about the transition density $p(\mathbf{x}_{\tau}\mid\mathbf{x}_{0})$ from across $\Omega$. Second, we use it to fix the stationary part of the learned transition density, boosting the sample efficiency when learning models of molecular dynamics. Third, it imposes an inductive bias of long-time dynamical behavior allowing for recovery of unbiased Boltzmann distribution for long-time horizons. Fourth, using BoPITO, we introduce a novel tunable sampling protocol interpolating ITO models trained on off-equilibrium simulation data and an unbiased equilibrium distribution.
@@ 3.2 Long-term dynamics inductive bias for ITO
$s_{\bm{\theta}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}},{\bm{\theta}}),$ (14)
where $s_{\mathrm{eq}}(\mathbf{x}^{t}_{\text{diff}},t_{\text{diff}})$ is the score of a pre-trained surrogate of the equilibrium distribution model, $10<\hat{\lambda}<$ 1 is a hyper-parameter and $s_{\mathrm{dyn}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}},{\bm{\theta}})$ accounts for the time-dependent components. As $N\rightarrow\infty$, $s_{\mathrm{eq}}$ dominates and the model samples from the equilibrium model, see Appendix A.2 for an example. By interpreting the score field as a velocity field, we can reformulate the DM as a continuous normalizing flow which we can use for Metropolized sampling of the unbiased equilibrium distribution (Klein et al., 2023). The proposed factorization is principled, corresponding to a separation of the score of the transition density into a stationary and dynamic part, where the first part is considered known, see Appendix A.3 for details. In practice, we first train $s_{\mathrm{eq}}$ (if not provided) using equilibrium data. Then we train $s_{\mathrm{dyn}}$ with unbiased, possibly off-equilibrium, MD data while keeping $s_{\mathrm{eq}}$ fixed.
@@ 3.3 BoPITO interpolators
A significant issue in MD simulations is the unbiased sampling of transitions over high free-energy barriers, e.g. channels in $\mu(\mathbf{x})$ with very little probability mass, typically coinciding with mixing on $\Omega$. This difficulty leads to an inability to predict time-correlation statistics for large $N\tau$. With BoPITO we can define an interpolation between models trained on off-equilibrium simulations and the equilibrium distribution. In this manner, we can approximate dynamics not seen explicitly in the data. For a given maximum training lag, $N_{\text{max}}$, we define a BoPITO interpolator as a model with the following score function:
$s_{\bm{\theta}}(\mathbf{x}^{t_{\text{diff}}},\mathbf{x}_{t},N_{\text{int}},t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N_{\text{int}}}s_{\mathrm{dyn}}(\mathbf{x}^{t_{\text{diff}}},\mathbf{x}_{t},N_{\text{max}},t_{\text{diff}}),$ (15)
@@
where $O_{N_{\text{int}}}$ is the dynamic observable estimated with samples from the BoPITO interpolator. This way BoPITO can integrate off-equilibrium and equilibrium MD with experimental data. In practice, we find the interpolator to generate some high-energy states. However, we can alleviate the high-energy structures by alternating long-lag interpolation steps with short-lag non-interpolation steps for local relaxation.

### v2

@@ Abstract
Accurate prediction of thermodynamic properties is essential in drug discovery and materials science. Molecular dynamics (MD) simulations provide a principled approach to this task, yet they typically rely on prohibitively long sequential simulations. Implicit Transfer Operator (ITO) Learning offers a promising approach to address this limitation by enabling stable simulation with time steps orders of magnitude larger than MD. However, to train ITOs, we need extensive, unbiased MD data, limiting the scope of this framework. Here, we introduce Boltzmann Priors for ITO (BoPITO) to enhance ITO learning in two ways. First, BoPITO enables more efficient data generation, and second, it embeds inductive biases for long-term dynamical behavior, simultaneously improving sample efficiency by one order of magnitude and guaranteeing asymptotically unbiased equilibrium statistics. Furthermore, we showcase the use of BoPITO in a new tunable sampling protocol interpolating between ITOs trained on off-equilibrium simulations and an equilibrium model by incorporating unbiased correlation functions. Code is available at https://github.com/olsson-group/bopito.
@@ 1 Introduction
Here, we introduce Boltzmann Priors for Implicit Transfer Operator (BoPITO) learning (Figure 1). BoPITO leverages pre-trained Boltzmann Generators (BG) (Noé et al., 2019; Viguera Diez et al., 2024; Klein & Noé, 2024; Köhler et al., 2023; 2020; Midgley et al., 2024; 2023) as priors to enable data-efficient training of ITO models, leading to one order of magnitude reduction in the simulation data needed for training. As the BG encodes the invariant measure or Boltzmann distribution of the dynamics encoded by the transition density, BoPITO, by construction, guarantees asymptotically unbiased equilibrium statistics. In this way, BoPITO can combine off-equilibrium data and biased data encoded into a BG prior to train ITO models that predict MD across multiple time-scales. Using BoPITO we introduce a new sampling strategy to recover approximate dynamics from biased off-equilibrium data, a BG prior, and unbiased time-correlation data, providing a new method for inverse problems for molecular systems.
@@ 2.1 Molecular dynamics and observables
Unfortunately, stable MD simulations rely on using time discretizations, $\tau$, on the order of $10^{-15}\,\mathrm{s}$, whereas properties such as protein folding or ligand unbinding occur on much longer time scales, $\sim 10^{-3}-10^{-1}\,\mathrm{s}$. Due to the sequential nature of these simulations, an impractical number of simulation steps are necessary to predict these properties in an unbiased fashion. As a result, most MD data will be ‘off-equilibrium’, e.g. trajectories exploring one or a few of the modes of the Boltzmann distribution $\mu(\mathbf{x})$. See Appendix A.1 for more precise definitions of configuration space, off-equilibrium, unbiased and biased MD.
@@ 2.2 Boltzmann Generators
Boltzmann Generators (BG) (Noé et al., 2019) are a generative machine learning approach to draw samples i.i.d. from the Boltzmann distribution, $\mu(\mathbf{x})$, of physical many-body systems (eq. 1). BGs learn a diffeomorphic map, $\mathcal{F}_{\bm{\theta}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$, from a latent space, equipped with a simple base density $p(\mathbf{z})$, to configurational space of a physical system, such that the push-forward density $\hat{p}(\mathbf{x})=\mathcal{F}_{\bm{\theta}}\#p(\mathbf{z})$ closely approximates the Boltzmann distribution (eq. 1). In practice, $\mathcal{F}_{\bm{\theta}}$ is implemented using an invertible neural network architecture with tractable Jacobian determinants to enable efficient sampling and exact sample likelihood computation (Chen et al., 2018; Papamakarios et al., 2021). BGs are trained either with approximately equilibrated simulation data or with biased simulations, from i.e. enhanced sampling, by employing appropriate reweighting (Ferrenberg & Swendsen, 1989; Shirts & Chodera, 2008). Unbiased samples can then be generated by importance re-sampling (Nicoli et al., 2020). Similarly, unbiased expectations can be computed using the importance weights $w(\mathbf{x})=e^{-\beta U(\mathbf{x})}/\hat{p}(\mathbf{x})$. As such, BGs are surrogates of the target Boltzmann distribution, but do not model time-correlation statistics.
@@ 2.3 Transfer operators
where $\lambda_{i}(\tau)$ are the eigenvalues, $\psi_{i}$ and $\phi_{i}$ are the corresponding right and left eigenfunctions, respectively, and $\langle f|g\rangle=\int_{\Omega}f(\mathbf{x})g(\mathbf{x})\,\mathrm{d}\mathbf{x}$. The eigenvalues $\lambda_{i}(\tau)$ depend on the parameter $\tau$ and are related to the characteristic relaxation rates $\kappa_{i}$ by $|\lambda_{i}(\tau)|=\exp(-\tau\kappa_{i})$. The right and left eigenfunctions are related by the stationary density, such that $\phi_{i}(\mathbf{x})=\mu(\mathbf{x})\psi_{i}(\mathbf{x})$. In reversible dynamics, all eigenvalues $\lambda_{i}$ are real and lie within the interval $-1<\lambda_{i}\leq 1$. Notably, there is one eigenvalue $\lambda_{1}=1$, with corresponding eigenfunctions $\psi_{1}(\mathbf{x})=\mathbf{1}$ and $\phi_{1}(\mathbf{x})=\mu(\mathbf{x})$.
@@ 2.4 Diffusion models
$\mathrm{d}\mathbf{x}^{t_{\mathrm{diff}}}=f(\mathbf{x}^{t_{\mathrm{diff}}},t_{\mathrm{diff}})\,\mathrm{d}t_{\mathrm{diff}}+g({t_{\mathrm{diff}}})\,\mathrm{d}W,$ (6)
where ($0<t_{\mathrm{diff}}<T$) is the diffusion time, $f$ and $g$ are chosen functions, and $\mathrm{d}W$ is a Wiener process. To generate samples from the data distribution, $p(x^{0})$, we can sample from $p(\mathbf{x}^{T})$ and solve the backward diffusion process (‘denoising process’) (Anderson, 1982),
$\mathrm{d}\mathbf{x}^{t_{\mathrm{diff}}}=\left[f(\mathbf{x}^{t_{\mathrm{diff}}},{t_{\mathrm{diff}}})-g^{2}({t_{\mathrm{diff}}})\nabla_{\mathbf{x}^{t_{\mathrm{diff}}}}\log p(\mathbf{x}^{t_{\mathrm{diff}}}|{t_{\mathrm{diff}}})\right]\mathrm{d}{t_{\mathrm{diff}}}+g({t_{\mathrm{diff}}})\,\mathrm{d}W.$ (7)
This backward process is approximated using a deep neural network $\nabla_{\mathbf{x}^{t_{\mathrm{diff}}}}\log p_{\bm{\theta}}(\mathbf{x}^{t_{\mathrm{diff}}}|{t_{\mathrm{diff}}})=s_{\bm{\theta}}(\mathbf{x}^{t_{\mathrm{diff}}},{t_{\mathrm{diff}}})$. The learned backward SDE can be recast into a probability flow ODE (Maoutsa et al., 2020; Song et al., 2021), which in turn can be interpreted as a continuous normalizing flow (Chen et al., 2018) which facilitates efficient sampling and enables sample likelihood evaluation.
@@ 2.5 Implicit Transfer Operator Learning
See Appendix A.3 for more details and pseudo-code on training and sampling algorithms.
3 Related work
Sampling the Boltzmann distribution
Latent space simulators and Coarse graining
Classical (fine-grained) molecular dynamics simulation can be coarse-grained beyond the Born-Oppenheimer approximation, such that atomic nuclei are merged together into ‘beads.’(Noid, 2023) This strategy, in principle enables faster simulations due to the smaller number of particles and the acceleration of kinetics caused by the coarse-graining (CG) operation (Nüske et al., 2019). CG models can be estimated to closely approximate the thermodynamics of the corresponding fine-grained system through the optimization of two equivalent variational bounds (Noid et al., 2008; Lyubartsev & Laaksonen, 1995; Ercolessi & Adams, 1994). A relaxation of this bound was recently proposed to also allow for probabilistic reconstruction of fine-grained configurations (Chennakesavalu et al., 2023). These bounds have been used extensively to build CG force-field models (Husic et al., 2020; Wang et al., 2019a; Majewski et al., 2023; Charron et al., 2023) including implicit solvation models (Chen et al., 2021; Katzberger & Riniker, 2024), using deep neural networks due to their ability to capture multibody terms (Wang et al., 2021) to ultimately accelerate the prediction of equilibrium properties of molecular systems. Similarly, the development of ‘latent space simulators’ where a learned, typically low-dimensional, latent space equipped either with a propagator (Sidky et al., 2020; Chennakesavalu et al., 2023; Wang et al., 2024), or not (Wang & Gómez-Bombarelli, 2019), is learned to enable efficient simulation. These approaches, in general aim to accelerate molecular simulations akin to BoPITO, yet due to the CG operation, the molecular dynamics (kinetics) will be accelerated, and detailed knowledge of the unbiased dynamics are needed to correct this (Nüske et al., 2019; Crommelin & Vanden-Eijnden, 2011). Concurrent work, presents MDGen where conformational states are tokenized and in turn used to generate multiple frames of a MD trajectory jointly (Jing et al., 2024).
Transfer Operator surrogates
Analysis of MD data often involves building transfer operator surrogates such as Markov state models (MSM) (Schütte et al., 1998; Prinz et al., 2011; Swope et al., 2004; Husic & Pande, 2018), time-lagged independent components analysis (Molgedey & Schuster, 1994; Ziehe & Müller, 1998; Pérez-Hernández et al., 2013), Markov field models or dynamic graphical models, (Olsson & Noé, 2019; Mardt et al., 2022; Hempel et al., 2022), VAMPnets (Mardt et al., 2018; Wu & Noé, 2019), or observable operator models (Wu et al., 2015). Markov state models are time-space discrete approximations of the transfer operator and Deep Generative MSM (Wu et al., 2019) and VAMPnets (Mardt et al., 2018) learn the space discretization through deep neural networks. Dynamic graphical models or Markov field models (Hempel et al., 2022) represent a time-space discrete approximation of the transfer operator injecting a (conditional) independence assumption of molecular subsystems, when modeling the transition probability, leading to better scaling for systems with poor time-scale separation (Olsson & Noé, 2019). Apart from ITO (Schreiner et al., 2023), several other deep generative approaches for modeling the transition density of molecular dynamics have recently proposed. Timewarp where a normalizing flow is used to encode the transition density (Klein et al., 2023) with limited transferability to enable metropolized sampling of unbiased equilibrium distributions (Hastings, 1970). Score dynamics use a DDPM to model the displacements of an initial configuration towards a time-lagged one, achieving picosecond time-steps simulation and limited transferability (Hsu et al., 2024). However, unlike in the context of MSMs (Trendelkamp-Schroer et al., 2015; Rosta & Hummer, 2014; Wu et al., 2014; 2016), there are no deep generative transition density surrogates available leveraging available information about the equilibrium distribution — BoPITO is one such method.
@@ 4 Boltzmann Priors for Implicit Transfer Operator Learning
BoPITO uses pre-trained models of the equilibrium distribution to improve ITO learning in four ways. First, it helps ensure broad sampling of $\Omega$ proportional to the equilibrium distribution for subsequent MD simulations yielding information about the transition density $p(\mathbf{x}_{\tau}\mid\mathbf{x}_{0})$ from across $\Omega$. Second, we use it to fix the stationary part of the learned transition density, boosting the sample efficiency when learning models of molecular dynamics. Third, it imposes an inductive bias of long-time dynamical behavior allowing for recovery of unbiased Boltzmann distribution for long-time horizons. Fourth, using BoPITO, we introduce a novel tunable sampling protocol interpolating ITO models trained on off-equilibrium simulation data and an unbiased equilibrium distribution.
@@ 4.2 Long-term dynamics inductive bias for ITO
$s(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}},\bm{\theta})=s_{\mathrm{eq}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N}s_{\mathrm{dyn}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}},{\bm{\theta}}),$ (14)
where $s_{\mathrm{eq}}(\mathbf{x}^{t}_{\text{diff}},t_{\text{diff}})$ is the score of a pre-trained surrogate of the equilibrium distribution model, $10<\hat{\lambda}<$ 1 is a hyper-parameter and $s_{\mathrm{dyn}}(\mathbf{x}_{t+N\tau}^{t_{\text{diff}}},\mathbf{x}_{t},N,t_{\text{diff}},{\bm{\theta}})$ accounts for the time-dependent components. As $N\rightarrow\infty$, $s_{\mathrm{eq}}$ dominates and the model samples from the equilibrium model, see Appendix A.4 for an example. By interpreting the score field as a velocity field, we can reformulate the DM as a continuous normalizing flow which we can use for Metropolized sampling of the unbiased equilibrium distribution (Klein et al., 2023). The proposed factorization is principled, corresponding to a separation of the score of the transition density into a stationary and dynamic part, where the first part is considered known, see Appendix A.5 for details. In practice, we first train $s_{\mathrm{eq}}$ (if not provided) using equilibrium data. Then we train $s_{\mathrm{dyn}}$ with unbiased, possibly off-equilibrium, MD data while keeping $s_{\mathrm{eq}}$ fixed and we choose $\hat{\lambda}$ performing hyper-parameter optimization, see Appendix A.6 for details. We discuss alternative formulations of the score in Appendix A.7.
@@ 4.3 BoPITO interpolators
A significant issue in MD simulations is the unbiased sampling of transitions over high free-energy barriers, e.g. channels in $\mu(\mathbf{x})$ with very little probability mass, typically coinciding with mixing on $\Omega$. This difficulty leads to an inability to predict time-correlation statistics for large $N\tau$. With BoPITO we can define an interpolation between models trained on off-equilibrium simulations and the equilibrium distribution. In this manner, we can approximate dynamics not seen explicitly in the data. For a time-dependent model, $s_{\mathrm{dyn}}$, trained on off-equilibrium simulations with maximum training lag, $N_{\text{max}}$, we define a BoPITO interpolator as a model with the following score function:
$s(\mathbf{x}^{t_{\text{diff}}},\mathbf{x}_{t},N_{\text{int}},t_{\text{diff}})=s_{\mathrm{eq}}(\mathbf{x}^{t_{\text{diff}}},t_{\text{diff}})+\hat{\lambda}^{N_{\text{int}}}s_{\mathrm{dyn}}(\mathbf{x}^{t_{\text{diff}}},\mathbf{x}_{t},N_{\text{max}},t_{\text{diff}}),$ (15)
@@
where $O_{N_{\text{int}}}$ is the dynamic observable estimated with samples from the BoPITO interpolator. This way BoPITO can integrate off-equilibrium and equilibrium MD with experimental data. In practice, we find the interpolator to generate some high-energy states. However, we can alleviate the high-energy structures by alternating long-lag interpolation steps with short-lag non-interpolation steps for local relaxation. We discuss how BoPITO interpolators alleviate out-of-distribution issues in Appendix A.8.